A chemist needs 160 milliliters of a 64% solution but has only 52% and 68% solutions available. Find how many milliliters of each that should be mixed to get the desired solution.
Answer by stanbon(75887)

A chemist needs 160 milliliters of a 64% solution but has only 52% and 68% solutions available. Find how many milliliters of each that should be mixed to get the desired solution.
-----
Equation:
active + active = active
---
0.52x + 0.68(160-x) = 0.64*160
-----
Multiply thru by 100 to get:
52x + 68*160 - 68x = 64*160
-16x = -4*160
x = 4*10
x = 40 ml (amt of 52% solution needed)
---
160-40 = 120 ml (amt of 68% solution needed)
=================================
Cheers,
Stan H.
==========